The friends send emails about the opening night to 128 people


They ask each person contacted to forward an email to five people on exchange for entry into a prize draw for a free meal for 2


Based on the assumptions below, how many people who received an email would you expect to attend the opening night.


Assumptions:

Half of the 128 people forward the email to 5 others


1/4 of these forward the email to 5 others


1/8 these forward the email to 5 others


1/16 these forward the email to 5 others


No one receives the email more than once.


10% of the people who receive an email go to the opening night.



1
Expert's answer
2021-02-25T01:33:17-0500

Since There are total of 128 people.


From the given assumption,

The intersection of the above assumption occur as


"\dfrac{1}{16}" of total people send email to five others and 10% who recieve the email attend the opening night.


Number of person who send email to 5 others= "\dfrac{1}{16}\times 128=8"


since 10% also attend the openimg night i.e. 13 people.


So People who recieve the email and attend the wedding night must be the intersection of the above two calculated people i.e. 8


hence The number of people Who attend the opening night and recieve the email is 8.